find two fractions equivalent to the given fractions:
1/5?

Found 2 solutions by stanbon, mananth:
Answer by stanbon(75887)

find two fractions equivalent to the given fractions:
1/5?
-------
(1/10) + (1/10) = 2/10 = 1/5
------------------
Cheers,
Stan H.

Answer by mananth(16946)